OpenStudy (anonymous):

Find exact value using half-angle identity cos7pi/12

OpenStudy (jdoe0001):

\(\bf \cfrac{7}{12}\cdot 2\implies \cfrac{7}{6}\qquad thus \\ \quad \\ cos\left(\cfrac{7\pi}{12}\right)\implies cos\left(\cfrac{{\color{blue}{ \frac{7\pi}{6}}}}{2}\right)=\pm \sqrt{\cfrac{1+cos\left({\color{blue}{ \frac{7\pi}{6}}}\right)}{2}}\)

OpenStudy (anonymous):

Is that the final answer? When input that in my calculator I get a decimal answer. But thank you for helping

OpenStudy (jdoe0001):

check your Unit Circle, what's the \(\bf cos\left(\frac{7\pi}{6}\right)\quad ?\)

OpenStudy (anonymous):

-sqrt3/2

OpenStudy (jdoe0001):

thus \(\bf cos\left(\cfrac{{\color{blue}{ \frac{7\pi}{6}}}}{2}\right)=\pm \sqrt{\cfrac{1+cos\left({\color{blue}{ \frac{7\pi}{6}}}\right)}{2}}\\ \quad \\\implies cos\left(\cfrac{{\color{blue}{ \frac{7\pi}{6}}}}{2}\right)=\pm \sqrt{\cfrac{1+\left({\color{blue}{ -\frac{\sqrt{3}}{2}}}\right)}{2}}\) and then simplify the radicand

OpenStudy (anonymous):

Yay! Ok that's what I have so far! But when I punch that in the calculator I still get a stupid decimal :((

OpenStudy (jdoe0001):

yes, the calculator will give you the decimal value so that's correct, if you want the decimal value if you just want the simplified version, you'd need to simplify the fractions in the root to make them just 1 fraction

OpenStudy (anonymous):

So do I multiply the bottom 2 to the top

OpenStudy (jdoe0001):

perform the SUM in the numerator, and keep in mind that \(\bf \cfrac{\quad \frac{a}{b}\quad }{\frac{c}{d}}\implies \cfrac{a}{b}\cdot \cfrac{d}{c}\)

OpenStudy (anonymous):

I'm sorry I'm a total airhead when it comes to math

OpenStudy (anonymous):

So multiply by the reciprocal :)

OpenStudy (jdoe0001):

yes, because \(\bf cos\left(\cfrac{ \frac{7\pi}{6}}{2}\right)=\pm \sqrt{\cfrac{1+\left({\color{blue}{ -\frac{\sqrt{3}}{2}}}\right)}{2}} \implies cos\left(\cfrac{ \frac{7\pi}{6}}{2}\right)=\pm \sqrt{\cfrac{1 -\frac{\sqrt{3}}{2}}{\frac{2}{1}}}\)

OpenStudy (jdoe0001):

so... what would you get for \(\bf 1 -\cfrac{\sqrt{3}}{2}\) ?

OpenStudy (anonymous):

Ok so far after that I have sqrt1+cos-sqrt3 and that's inside a big sqrt

OpenStudy (jdoe0001):

hmmm recall your fractions additions \(\bf cos\left(\cfrac{ \frac{7\pi}{6}}{2}\right)=\pm \sqrt{\cfrac{1+\left({\color{blue}{ -\frac{\sqrt{3}}{2}}}\right)}{2}} \implies cos\left(\cfrac{ \frac{7\pi}{6}}{2}\right)=\pm \sqrt{\cfrac{1 -\frac{\sqrt{3}}{2}}{\frac{2}{1}}} \\ \quad \\ cos\left(\cfrac{ \frac{7\pi}{6}}{2}\right)=\pm \sqrt{\cfrac{\frac{2-\sqrt{3}}{2}}{\frac{2}{1}}}\)

OpenStudy (jdoe0001):

so you'd have two fractions, so recall \(\bf \cfrac{\quad \frac{a}{b}\quad }{\frac{c}{d}}\implies \cfrac{a}{b}\cdot \cfrac{d}{c}\)

OpenStudy (anonymous):

Doesn't the bottom cancel out?

OpenStudy (jdoe0001):

hmmmm flip the denominator, let's see if it does cancel out

OpenStudy (anonymous):

Not if I flip it because then it's 1/2 :/

OpenStudy (jdoe0001):

\(\bf \cfrac{\frac{2-\sqrt{3}}{2}}{\frac{2}{1}}\implies \cfrac{2-\sqrt{3}}{2}\cdot \cfrac{1}{2}\)

OpenStudy (anonymous):

Omg I forgot the two before the -sqrt3

OpenStudy (jdoe0001):

\(\bf cos\left(\cfrac{ \frac{7\pi}{6}}{2}\right)=\pm \sqrt{\cfrac{1+\left({\color{blue}{ -\frac{\sqrt{3}}{2}}}\right)}{2}} \implies cos\left(\cfrac{ \frac{7\pi}{6}}{2}\right)=\pm \sqrt{\cfrac{1 -\frac{\sqrt{3}}{2}}{\frac{2}{1}}} \\ \quad \\ cos\left(\cfrac{ \frac{7\pi}{6}}{2}\right)=\pm \sqrt{\cfrac{\frac{2-\sqrt{3}}{2}}{\frac{2}{1}}}\implies cos\left(\cfrac{ \frac{7\pi}{6}}{2}\right)=\pm\sqrt{\cfrac{2-\sqrt{3}}{4}}\)

OpenStudy (anonymous):

That looks so much more clear. Thank you so so so much !!!! :)

OpenStudy (jdoe0001):

yw